# A Tutorial on Bayesian Analysis of Linear Shock Compression Data

### 06 Mar 2026 v1

@@ 3 Bayesian Analysis
The interpretation of $p(\beta|\sigma^{2},Y)$ is that if $\sigma^{2}$ is known, then the posterior distribution of $\beta$ is normal. This hierarchical factorization of the joint posterior distribution suggests that one way to sample a $\beta$ from its marginal posterior distribution is to first sample $\sigma^{2}\sim IG(a,b)$ following Eq. (12) and then to sample $\beta\sim N(\hat{\beta},\sigma^{2}(X^{\prime}X)^{-1})$} following Eq. (11). Since the samples of $\beta$ and $\sigma^{2}$ are from the joint posterior distribution of $(\beta,\sigma^{2})$, they can be used for simulating synthetic data or for estimating integrals using Monte Carlo. For example, the probability that $C_{0}$ is greater than 4 can be estimated as
$N^{-1}\sum_{i=1}^{N}I\{C_{0}^{(i)}>4\},$ (13)
where $\beta^{(i)}=(C_{0}^{(i)},S^{(i)})^{\prime}$ is the $i$th of $N$ samples of $\beta$ and $I\{C_{0}^{(i)}>4\}$ is an indicator function that equals one if $C_{0}^{(i)}>4$ and equals zero otherwise.
@@ 6 Discussion
In general, posterior distributions can be sampled from approximately or computed numerically when they cannot be sampled directly or are not analytically available. This would occur for our regression model if the prior distribution was not a normal-inverse-gamma distribution, for example. One approach is to use an iterative, probabilistic algorithm called Markov Chain Monte Carlo (MCMC) to obtain approximate samples from the posterior distribution (Gelman et al., 1995, Ch. 11). The emcee package in Python is a popular choice for implementing MCMC (Foreman-Mackey et al., 2013), though care should be taken if the number of parameters to infer is large (Huijser et al., 2022). An MCMC algorithm called a Gibbs sampler can also be used to draw samples from the posterior distribution of the linear model parameters, as described in Chib (2001, Sec. 8.3). Note that these MCMC algorithms produce a sequence of correlated samples from the posterior distribution, whereas the direct sampling methods presented in this paper produce independent samples. Alternatively, since the linear regression model examined in this paper only has three parameters, the posterior distribution can be evaluated on a three-dimensional grid. That is, the prior distribution is multiplied by the likelihood for each grid cell to obtain the non-normalized posterior distribution, and then these values are normalized to sum to one to approximate the actual posterior distribution.
@@ Acknowledgements
$f(x)\propto\left[1+\frac{1}{\nu}(x-\mu)^{\prime}\Sigma^{-1}(x-\mu)\right]^{-(\nu+p)/2},$ (A1)
for $x\in\mathbb{R}^{p}$, which is notated
@@
The proportionality constant is a normalizing constant that ensures $f(x)$ integrates to one over $\mathbb{R}^{p}$. As $\nu\rightarrow\infty$, the distribution converges to a normal distribution with mean $\mu$ and covariance matrix $\Sigma$ since
$\lim_{\nu\rightarrow\infty}f(x)\propto\exp\left(-\frac{1}{2}(x-\mu)^{\prime}\Sigma^{-1}(x-\mu)\right),$ (A3)
@@
$\lim_{n\rightarrow\infty}\left(1+\frac{z}{n}\right)^{n}=\exp(z)$ (A4)
with $n=\nu/2$ and $z$ given by the quadratic form in Eq. (A1) divided by two. When $\nu=1$, the $t$-distribution is a Cauchy distribution.
@@
Figure 10: The $t$-distribution converges to a normal distribution as the degrees of freedom increases to infinity. For this plot, the dimension of the $t$-distribution is $p=1$, the mean is $\mu=0$ and the scale matrix is $\Sigma=1.5$, which is a scalar.
@@
$s^{2}=Y^{\prime}(I_{n}-X(X^{\prime}X)^{-1}X^{\prime})Y.$ (B8)

### 15 Mar 2026 v2

@@ 3 Bayesian Analysis
The interpretation of $p(\beta|\sigma^{2},Y)$ is that if $\sigma^{2}$ is known, then the posterior distribution of $\beta$ is normal. This hierarchical factorization of the joint posterior distribution suggests that one way to sample a $\beta$ and $\sigma^{2}$ from their joint posterior distribution is to first sample $\sigma^{2}\sim IG(a,b)$ following Eq. (12) and then to sample $\beta\sim N(\hat{\beta},\sigma^{2}(X^{\prime}X)^{-1})$} following Eq. (11). Such samples of $\beta$ and $\sigma^{2}$ can be used for simulating synthetic data or for estimating integrals using Monte Carlo. For example, a Monte Carlo estimate of the probability that $C_{0}$ is greater than 4, conditioned on the available data, is
$p(C_{0}>4|Y)\approx N^{-1}\sum_{i=1}^{N}I\{C_{0}^{(i)}>4\},$ (13)
where $\beta^{(i)}=(C_{0}^{(i)},S^{(i)})^{\prime}$ is the $i$th of $N$ samples of $\beta$ and $I\{\cdot\}$ is an indicator function.
@@ 6 Discussion
In general, posterior distributions can be sampled from approximately or computed numerically when they cannot be sampled directly or are not analytically available. This would occur for our regression model if the prior distribution was not a normal-inverse-gamma distribution, for example. One approach to obtaining approximate samples from a posterior distribution is to use an iterative, probabilistic algorithm called Markov Chain Monte Carlo (MCMC) (Gelman et al., 1995, Ch. 11). The emcee package in Python is a popular choice for implementing MCMC (Foreman-Mackey et al., 2013), though care should be taken if the number of parameters to sample is large (Huijser et al., 2022). An MCMC algorithm called a Gibbs sampler can also be used to draw samples from the posterior distribution of the linear model parameters, as described in Chib (2001, Sec. 8.3). Note that these MCMC algorithms produce a sequence of correlated samples from the posterior distribution, whereas the direct sampling methods presented in this paper produce independent samples. Alternatively, since the linear regression model examined in this paper only has three parameters, the posterior distribution can be evaluated on a three-dimensional grid. That is, the prior distribution is multiplied by the likelihood for each grid cell to obtain the non-normalized posterior distribution, and then these values are normalized to sum to one to approximate the actual posterior distribution.
@@ Acknowledgements
$f(x)=C(\nu,p,\Sigma)\left[1+\frac{1}{\nu}(x-\mu)^{\prime}\Sigma^{-1}(x-\mu)\right]^{-(\nu+p)/2},$ (A1)
for $x\in\mathbb{R}^{p}$ and
$C(\nu,p,\Sigma)=\frac{\Gamma((\nu+p)/2)}{\Gamma(\nu/2)\nu^{p/2}\pi^{p/2}|\Sigma|^{1/2}},$ (A2)
which is notated
@@
The normalizing constant, $C(\nu,p,\Sigma)$, ensures that $f(x)$ integrates to one over $\mathbb{R}^{p}$. When $\nu=1$, the multivariate $t$-distribution is a multivariate Cauchy distribution.
We show that the multivariate $t$-distribution converges to the multivariate normal distribution as $\nu\rightarrow\infty$ in two steps. First, it follows from Abramowitz and Stegun (1948, eqn. 6.1.46) that
$\lim_{\nu\rightarrow\infty}C(\nu,p,\Sigma)=\frac{1}{2^{p/2}\pi^{p/2}|\Sigma|^{1/2}}.$ (A4)
Second, the limit of the second term in the probability density function in Eq. (A1) is
$\lim_{\nu\rightarrow\infty}\left[1+\frac{1}{\nu}(x-\mu)^{\prime}\Sigma^{-1}(x-\mu)\right]^{-(\nu+p)/2}=\exp\left(-\frac{1}{2}(x-\mu)^{\prime}\Sigma^{-1}(x-\mu)\right),$ (A5)
@@
$\lim_{\nu\rightarrow\infty}\left(1+\frac{z}{\nu}\right)^{-\nu/2}=\exp\left(-\frac{z}{2}\right)$ (A6)
with $z=(x-\mu)^{\prime}\Sigma^{-1}(x-\mu)$. Hence, combining the two limits, it follows that the density function of the multivariate $t$-distribution converges to the density function of a multivariate normal distribution with mean $\mu$ and covariance matrix $\Sigma$. The first step can be simplified when $p=2$ since the numerator of $C(\nu,p,\Sigma)$ reduces to $(\nu/2)\Gamma(\nu/2)$ by Abramowitz and Stegun (1948, eqn. 6.1.15), which implies that the normalizing constant does not depend on $\nu$.
@@
Figure 10: The $t$-distribution converges to a normal distribution as the degrees of freedom increases to infinity. For this plot, the dimension of the $t$-distribution is $p=1$, the mean is $\mu=0$, and the scale matrix is $\Sigma=1.5$, which is a scalar.
@@
$\nu s^{2}=Y^{\prime}(I_{n}-X(X^{\prime}X)^{-1}X^{\prime})Y.$ (B8)